find the dimensions of a rectangle with perimeter 36 square units and area 65 square units.whitch of the following could be the dimensions of the rectangles?can you please explain this to me?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

find the dimensions of a rectangle with perimeter 36 square units and area 65 square units.which of the following could be the dimensions of the rectangles?can you please explain this to me?
-------------------------------------------
Perimeter = 2(L+W)=36 units
Area = LW = 65 sq. units
----------------------------
L+W = 18
LW = 65
-----------------
L = 18-W
-----
Substitute for "L":
(18-W)W = 65
-W^2 + 18W - 65 = 0
W^2 - 18W + 65 = 0
---
Factor:
(W-13)(W-5) = 0
W = 13 or W = 5 units
If Width = 13, Length = 5 units
=====================================
Cheers,
Stan H.

Answer by mananth(16949)

perimeter 36
area 65
2L+2W= 36 ( perimeter )
L*W = 65 ( area)
L= 65/w
.
2L+2W= 36
2* 65 /W + 2W= 36
130+2W^2 / W = 36
130 +2W^2 =36W
2W^2-36W+130=0
2(W^2-18W+65)=0
2(w-13)(W-5)=0
W= 13 Or 5
The dimensions are 13 by 5 units